# Kinetic Interacting Particle Langevin Monte Carlo

### v2

@@ 5.2 Wisconsin Cancer Data
We follow again an experimental procedure that is similar to the one outlined in [47] and [3] and make comparisons between the algorithms on a more realistic dataset: the Wisconsin Cancer Data. Again, we use the logistic regression LVM model, outlined above. This task is a binary classification, to determine from 9 features gathered from tumors and 693 data points labelled as either benign or malignant. The latent variables correspond to the features extracted from the data. The task in this case is to seek to model the behaviour as accurately as possible through the logistic regression LVM.
@@
Figure 3: Wisconsin Dataset. The performance of MPGD, MPGDnc, KIPLMC1 and KIPLMC2 algorithms are compared on a logistic regression experiment for the Wisconsin Cancer Dataset. In (a), we show the behaviour of the $\theta_{n}$ iterates for the small step-size of $\eta=0.01$, where all algorithms converge as desired. (b) shows the behaviour $\theta_{n}$ iterates for a step-size where some algorithms explode. In (c) we compare the distributions of the algorithms over different step-sizes. MPGDnc is the MPGD algorithm without gradient correction. The particle number, is chosen to be $N=50$ and $\gamma=1.2$.
Most notably in this experiment, we can observe in Fig. 3 the importance of the discretisation KIPLMC2. In particular, this algorithm displays great stability w.r.t. the choice of the step-size. For small values of the step-size, the MPGDnc algorithm performs with the lowest variance in all cases where it converges, until it explodes where step-sizes become too large (for more detail see C.2). This is remedied by implementing a gradient correction trick, discussed in [49], though we show how the algorithm performs without for a comparison with KIPLMC1. Again, we note that the KIPLMC1 algorithm exhibits more variance in the $\theta$ estimation than the KIPLMC2 and MPGD algorithms. This is typically an advantage when working with convex problems, but this “sticky” behaviour might prove detrimental in the non-convex case [38, 3]. The injection of noise into the parameter estimation may help the method to escape local minima [1]. In these algorithms one can observe the strength of the KIPLMC2 algorithm, with fast convergence and smoother paths. The latter observation may be related to the differing orders of the algorithms in [53] and [25].
@@ 5.3 Bayesian Neural Network
Similarly to [47, 49] we also consider a Bayesian Neural Network (BNN) example to perform character classification on the MNIST dataset. This dataset and model provide a more challenging problem for the 1 and 2 algorithms as the posteriors are known to be multi-modal. MNIST contains 70’000 $28\times 28$ grey-scale images $\{f_{i}\}_{i=1}^{70000}\subset\mathbb{R}^{784}$. However, to avoid issues of high dimensionality, we consider a normalised subset of 1’000 characters containing only images of fours and nines, whose similarity should pose a challenge.
Following [47] we employ a two layer BNN with $\tanh$ activation function, softmax output layer and 40 dimensional latent space. i.e. we consider the probability of the data labels $l$ to be conditionally independent, given the features $f$ and the weights $x=(w,v)$, with model,
@@
$p_{\theta}(x,\{f_{i}\}_{i=1}^{1000})=\mathcal{N}(w;0,e^{2\alpha}\mathbb{I})\mathcal{N}(v;0,e^{2\beta}\mathbb{I})\prod_{i=1}^{1000}p(l|f_{i},x).$
Notice again that we are unable to confirm the validity of our assumptions. We will employ autograd methods from the JAX library to compute the gradients.
Figure 4: Bayesian Neural Network. (a) LPPD estimates as a function of step count, and (b) the relative error (for a discussion of these metrics see C.3). All algorithms hover around an relative error of 0.02 when converged. Shown are the averaged behaviour over 10 simulations, which were all run with $N=100$, $\gamma=1.9$ and $\eta=0.015$.
In Fig. 4 we make a comparison between the different algorithms over one run, comparing the log predictive point-wise density (LPPD), relative error and the density of the last estimate of $w$. The LPPD is the average log-likelihood assigned to the correct response, whilst the relative error is given as the accuracy of prediction on a test set (see C.3 for more detail). Observe the improved behaviour of the KIPLMC2 algorithm in LPPD (a), comparable behaviour in relative error (b). Note however that the KIPLMC2 algorithm converges to a different solution than all other algorithms in our simulations.
@@ 6 Conclusions
This paper extends a line of work on interacting particle, and more generally diffusion-based, algorithms for maximum marginal likelihood estimation. This paper has focused on considering alternatives to the IPSs proposed by [47, 3, 49] for the MMLE problem by considering an accelerated variant. We have shown that we can leverage the existing literature on underdamped Langevin diffusion for sampling [25, 20, 53, 51] to produce two algorithms with greater stability, added smoothness and exponential convergence, which concentrate onto the MMLE with quantitative nonasymptotic bounds on the estimation error as measured by ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}$. In particular, we show that the KIPLMC1 does attain an accelerated error rate for attaining an error of order $\varepsilon$ after $\widetilde{\mathcal{O}}(\sqrt{d_{x}}\varepsilon^{-1})$ steps, compared to the IPLA and PGD which require $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps.
Proposed algorithms within this manuscript are both based on a momentum-based diffusion - which can both be seen as a relaxation of the friction of the IPLA in [3] or as a noised version of the MPGD in [49]. Empirically, the algorithms perform well compared to the MPGD (with equal choice of step-sizes and friction coefficients). Specifically, the induced noise should improve guarantees in the non-convex setting, as discussed in [3, 57, 1]. The results presented here hold under assumptions of gradient-Lipschitzness and strong log-concavity, however, following the work done in the analysis of Langevin diffusions under nonconvexity and without global Lipschitz continuity (see, e.g., [71, 1, 44]), it is possible to extend these results to these cases, which is left for future work. Another interesting direction is to consider the setting in [36], using an accelerated algorithm for the high-dimensional linear models. Our methods can also be adapted to solving linear inverse problems as done by [5]. Similar ideas can be considered within the setting of Stein variational gradient descent as done in [63].
@@ Appendix A Preliminary results
Lemma A.1 (KIPLD as an underdamped Langevin Diffusion).
The KIPLD diffusion can be equivalently written as a single $d_{\theta}+Nd_{x}$-dimensional underdamped Langevin diffusion given by
${\mathrm{d}}{\mathbf{Z}}_{t}$ $=\bar{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t,$ (15)
${\mathrm{d}}\bar{{\mathbf{V}}}_{t}^{z}$ $=-{\gamma}\bar{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t-\nabla_{z}\bar{U}_{N}({\mathbf{Z}}_{t}){\mathrm{d}}t+\sqrt{\frac{2{\gamma}}{N}}{\mathrm{d}}{\mathbf{B}}_{t},$
where ${\mathbf{Z}}_{t}^{i}=N^{-1/2}{\mathbf{X}}_{t}^{i}$ for $i=1,\ldots,N$, ${\mathbf{Z}}_{t}=(\bm{\theta}_{t},{\mathbf{Z}}_{t}^{1},\ldots,{\mathbf{Z}}_{t}^{N})^{\top}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}}$, $\bar{{\mathbf{V}}}_{t}^{z_{i}}=N^{-1/2}{\mathbf{V}}_{t}^{x_{i}}$ and $\bar{{\mathbf{V}}}_{t}^{z}=({\mathbf{V}}_{t}^{\theta},\bar{{\mathbf{V}}}_{t}^{z_{1}},\ldots,\bar{{\mathbf{V}}}_{t}^{z_{N}})^{\top}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}}$, $({\mathbf{B}}_{t})_{t\geq 0}$ is a ${\mathbb{R}}^{d_{\theta}+Nd_{x}}$-valued Brownian motion, and
$\bar{U}_{N}(\theta,z_{1},\ldots,z_{N}):=\frac{1}{N}\sum_{i=1}^{N}U(\theta,\sqrt{N}z_{i}).$
@@
Recall the KIPLD system:
$\mathrm{d}\bm{\theta}_{t}$ $={\mathbf{V}}^{\theta}_{t}\mathrm{d}t$
$\mathrm{d}{\mathbf{X}}^{i}_{t}$ $={\mathbf{V}}^{x_{i}}_{t}\mathrm{d}t,\qquad\text{for }i=1,\ldots,N$
$\mathrm{d}{\mathbf{V}}^{x_{i}}_{t}$ $=-\gamma{\mathbf{V}}_{t}^{x_{i}}\mathrm{d}t-\nabla_{x}U(\bm{\theta}_{t},{\mathbf{X}}_{t}^{i})\mathrm{d}t+\sqrt{2\gamma}\mathrm{d}{\mathbf{B}}_{t}^{i},\qquad\text{for }i\in[N]$
We note now that, below, we generically use the gradient functions $\nabla_{\theta}U(\theta,x):\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{\theta}}$ and $\nabla_{x}U(\theta,x):\mathbb{R}^{d_{\theta}}\times\mathbb{R}^{d_{x}}\to\mathbb{R}^{d_{x}}$ even though their arguments may change. In other words, below $\nabla_{\theta}U(\theta,x)$ denotes the gradient of $U$ with respect to its first argument and $\nabla_{x}U(\theta,x)$ denotes the gradient of $U$ with respect to its second argument.
Let us introduce the following function:
$U_{N}(\theta,x)=U(\theta,\sqrt{N}x),$
where we note that the (full) gradient is given as,
$\nabla U_{N}(\theta,x)=\begin{pmatrix}\nabla_{\theta}U(\theta,\sqrt{N}x)\\ N^{-1/2}\nabla_{x}U(\theta,\sqrt{N}x)\end{pmatrix}.$
Let us now introduce $z_{\theta}\in\mathbb{R}^{d_{\theta}}$ and $z_{x}=(z_{1},\dots,z_{N})^{T}$ where $z_{i}\in\mathbb{R}^{d_{x}}$ for $i=1,\ldots,N$. We define
$\bar{U}_{N}(z_{\theta},z_{x})=\frac{1}{N}\sum_{i=1}^{N}U_{N}(z_{\theta},z_{i}).$ (16)
The gradients are given as
$\nabla_{z_{\theta}}\bar{U}_{N}(z_{\theta},z_{x})=\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(z_{\theta},\sqrt{N}z_{x}^{i}),\qquad\nabla_{z_{x}}\bar{U}_{N}(z_{\theta},z_{x})=\begin{pmatrix}N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{1})\\ \vdots\\ N^{-1/2}\nabla_{x}U(z_{\theta},\sqrt{N}z_{N})\end{pmatrix}.$ (17)
Next, let ${\mathbf{Z}}_{t}^{i}=N^{-1/2}{\mathbf{X}}_{t}^{i}$ for $i=1,\ldots,N$, ${\mathbf{Z}}_{t}=(\bm{\theta}_{t},{\mathbf{Z}}_{t}^{1},\ldots,{\mathbf{Z}}_{t}^{N})^{\top}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}}$, $\bar{{\mathbf{V}}}_{t}^{z_{i}}=N^{-1/2}{\mathbf{V}}_{t}^{x_{i}}$ and $\bar{{\mathbf{V}}}_{t}^{z}=({\mathbf{V}}_{t}^{\theta},\bar{{\mathbf{V}}}_{t}^{z_{1}},\ldots,\bar{{\mathbf{V}}}_{t}^{z_{N}})^{\top}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}}$. We can rewrite the system KIPLD as
${\mathrm{d}}{\mathbf{Z}}_{t}$ $=\bar{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t,$ (18)
@@
${\mathrm{d}}{{\mathbf{V}}}^{\theta}_{t}$ $=-\gamma{{\mathbf{V}}}^{\theta}_{t}{\mathrm{d}}t-\frac{1}{N}\sum_{i=1}^{N}\nabla_{\theta}U(\bm{\theta}_{t},\sqrt{N}{\mathbf{Z}}_{t}^{i}){\mathrm{d}}t+\sqrt{\frac{2\gamma}{N}}{\mathrm{d}}{\mathbf{B}}_{t}^{0},$
${\mathrm{d}}\bar{{\mathbf{V}}}_{t}^{z_{i}}$ $=-\gamma\bar{{\mathbf{V}}}_{t}^{z_{i}}{\mathrm{d}}t-N^{-1/2}\nabla_{x}U(\bm{\theta}_{t},\sqrt{N}{\mathbf{Z}}_{t}^{i}){\mathrm{d}}t+\sqrt{\frac{2\gamma}{N}}{\mathrm{d}}{\mathbf{B}}_{t}^{i},\qquad\text{for }i\in[N].$
where ${\mathbf{B}}^{0}_{t}$ is a ${\mathbb{R}}^{d_{\theta}}$-valued Brownian motion and ${\mathbf{B}}^{i}_{t}$ for $i=1,\ldots,N$ are ${\mathbb{R}}^{d_{x}}$-valued Brownian motions. Rewriting the system (18) in terms of $\bar{U}_{N}$, using (17), we obtain the result. $\Box$
We can also write the KIPLD in the following alternative way which will ease the analysis.
Lemma A.2 (KIPLD as a standard underdamped Langevin Diffusion).
The KIPLD diffusion can also be written as a single $d_{\theta}+Nd_{x}$-dimensional underdamped Langevin diffusion given by
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\widetilde{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t-N\nabla_{z}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{{2{\widetilde{\gamma}}}}{\mathrm{d}}{\mathbf{B}}_{t},$
where $\widetilde{{\mathbf{Z}}}_{t},\widetilde{{\mathbf{V}}}_{t}^{z}\in\mathbb{R}^{d_{x}+Nd_{\theta}}$ are given as,
$\widetilde{{\mathbf{Z}}}_{t}=(\bm{\theta}_{\sqrt{N}t},N^{-1/2}{\mathbf{X}}^{1}_{\sqrt{N}t},\dots N^{-1/2}{\mathbf{X}}^{N}_{\sqrt{N}t})^{\intercal},\qquad\widetilde{{\mathbf{V}}}_{t}^{z}=(\sqrt{N}{\mathbf{V}}^{\theta}_{\sqrt{N}t},{\mathbf{V}}^{x_{1}}_{\sqrt{N}t},\dots,{\mathbf{V}}^{x_{N}}_{\sqrt{N}t})^{\intercal},$
$({\mathbf{B}}_{t})_{t\geq 0}$ is a ${\mathbb{R}}^{d_{\theta}+Nd_{x}}$-valued Brownian motion, $\tilde{\gamma}=\sqrt{N}\gamma$ and
$\bar{U}_{N}(\theta,z_{1},\ldots,z_{N}):=\frac{1}{N}\sum_{i=1}^{N}U(\theta,z_{i}).$
Proof.
We use [25, Lemma 1] to prove this result. Consider the variables $\widetilde{{\mathbf{Z}}}_{t}$ and $\widetilde{{\mathbf{V}}}_{t}^{z}$. It is straightforward to observe that
${\mathrm{d}}\widetilde{{\mathbf{Z}}}_{t}=\widetilde{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t.$
Indeed the space rescaling accounts for the time-rescaling in the dynamics of $\widetilde{{\mathbf{Z}}}_{t}$. Now observe that
@@
${\mathrm{d}}\widetilde{{\mathbf{V}}}^{\theta}_{t}$ $=-N\gamma{\mathbf{V}}^{\theta}_{\sqrt{N}t}{\mathrm{d}}t-\nabla_{\theta}\sum_{i=1}^{N}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}{\mathrm{d}}{\mathbf{B}}_{t}^{0}$
$=-\sqrt{N}\gamma\widetilde{{\mathbf{V}}}_{t}^{\theta}{\mathrm{d}}t-N\nabla_{\theta}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}{\mathrm{d}}{\mathbf{B}}_{t}^{0}.$
Similarly we obtain for $i\in[N]$,
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{x_{i}}$ $=-\sqrt{N}\gamma{\mathbf{V}}^{i}_{\sqrt{N}t}{\mathrm{d}}t-\nabla_{x}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}{\mathrm{d}}B_{t}^{i}$
$=-\sqrt{N}\gamma\widetilde{{\mathbf{V}}}_{t}^{x_{i}}{\mathrm{d}}t-\nabla_{x}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}{\mathrm{d}}B_{t}^{i},$
where we recall our definition of the gradient of $\bar{U}_{N}$ from (17) to obtain the correct scaling in front of $\nabla_{x}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})$. Now, taking $\tilde{\gamma}=\sqrt{N}\gamma$, the result follows. $\Box$
Recall that the algorithms proposed seek to target with the $n$th step the analytic solution at time $n\eta$, where $\eta$ is the step-size of the algorithm for the rescaled system. Hence, by re-scaling $\tilde{\eta}=\frac{\eta}{\sqrt{N}}$, we can use algorithms for the rescaled system to produce estimates for KIPLMC1 and KIPLMC2. This is discussed below in B.4 and B.5. Given these results, it is natural to explore the properties of the function $\bar{U}_{N}$, as this new potential plays a crucial role in both rescalings. We note below some properties of $\bar{U}_{N}$ which will be useful in the proofs.
@@
The measure $\tilde{\pi}$, left stationary by the dynamics of (19), is given as
@@
This result follows directly by observing the rescaling given in Lemma A.2, both in the parameters and equation. For a discussion of the stationary measure of a standard underdamped Langevin diffusion see [55, Chapter 6].

### v3

@@ 5.2 Wisconsin Cancer Data
We follow again an experimental procedure that is similar to the one outlined in Kuntz et al. [35] and Akyildiz et al. [1] and make comparisons between the algorithms on a more realistic dataset: the Wisconsin Cancer Data. Again, we use the logistic regression LVM model, outlined above. This task is a binary classification, to determine from 9 features gathered from tumors and 693 data points labelled as either benign or malignant. The latent variables correspond to the features extracted from the data. The task in this case is to seek to model the behaviour as accurately as possible through the logistic regression LVM.
@@
Figure 4: MNIST Dataset. The performance of PGD, IPLA, MPGD, KIPLMC1 and KIPLMC2 algorithms are compared on a classification experiment for the MNIST Dataset. (a) shows the Log Pointwise Predictive Density, the average log probability assigned to the correct response. (b) shows the percentage error and (c) the posterior density of the weights $w$. For this experiment $N=100$, $\gamma=2$ and $\eta=0.01$.
Most notably in this experiment, we can observe in Fig. 3 the importance of the added stability that the momentum-based algorithms exhibit, with the use of more stable numerical integrators than Euler-Maruyama. In particular, these algorithms displays great stability w.r.t. the choice of the step-size, as well as, the acceleration one would expect. For small step-sizes, the PGD algorithm performs with low variance in all cases where it converges, until it explodes where step-sizes become too large . It is interesting to note however that in the large step-size regime, the momentum-based algorithms exhibit much greater momentum than in the smaller step-sizes. However, again, we note that the KIPLMC1 and KIPLMC2 algorithms exhibit more variance in the $\theta$ estimation than the PGD and MPGD algorithms, though it is damped when compared to the IPLA algorithm, due to the higher regularity of the solutions to KIPLD. This is typically an advantage when working with convex problems, but this “sticky” behaviour might prove detrimental in the non-convex case [27, 1]. The injection of noise into the parameter estimation may help the method to escape local minima [3].
@@ 5.3 Bayesian Neural Network
Similarly to Kuntz et al. [35], Lim et al. [37] we also consider a Bayesian Neural Network (BNN) example to perform character classification on the MNIST dataset. In this case our approach is similar to “cold” posterior estimation, as our sampling density is proportional to a density raised to the inverse temperature, as in Wenzel et al. [51].
This dataset and model provide a more challenging problem for the KIPLMC1 and KIPLMC2 algorithms as the posteriors are known to be multi-modal and hence do not satisfy assumption H1. MNIST contains 70’000 $28\times 28$ grey-scale images $\{f_{i}\}_{i=1}^{70000}\subset\mathbb{R}^{784}$. However, to avoid issues of high dimensionality, we consider a normalised subset of 1,000 characters containing only images of fours and nines, whose similarity should pose a challenge. Due to this smaller sample, we are also able to avoid the need for sub-sampling.
Following Kuntz et al. [35] we employ a two layer BNN with $\tanh$ activation function, softmax output layer and 40 dimensional latent space. i.e. we consider the probability of the data labels $l$ to be conditionally independent, given the features $f$ and the weights $x=(w,v)$, with model,
@@
$p_{\theta}(x,\{f_{i}\}_{i=1}^{1000})=$ $\mathcal{N}(w;0,e^{2\alpha})\mathcal{N}(v;0,e^{2\beta})\prod_{i=1}^{1000}p(l|f_{i},x).$
We will employ autograd methods from the PyTorch library to compute the gradients.
In Fig. 4 we make a comparison between the different algorithms over 100 runs, comparing the log predictive point-wise density (LPPD) and relative error. The LPPD is the average log-likelihood assigned to the correct response, whilst the relative error is given as the accuracy of prediction on a test set (see D.4 for more detail). Observe the improved behaviour of the momentum-based algorithms in LPPD, as well as relative error. We also note the significant acceleration observed for KIPLMC2, highlighting the effect of the distinct numerical integrators. Similarly, in Fig. 4 (c) we can see the more concentrated posterior distributions of the momentum-based algorithms. Thus it is clear that the KIPLMC1 and KIPLMC2 algorithms perform competitively against other particle-based methods, even in very complex problems, in which our assumptions may not be satisfied.
@@ 6 Conclusions
This paper extends a line of work on interacting particle-, and more generally, diffusion-based algorithms for maximum marginal likelihood estimation. This paper has focused on alternatives to the interacting particle systems (IPSs) proposed by Kuntz et al. [35], Akyildiz et al. [1], Lim et al. [37] for the MMLE problem by considering an accelerated variant. We have shown that we can leverage the existing literature on underdamped Langevin diffusion for sampling [19, 16, 41, 39] to produce two algorithms with greater stability, added smoothness and exponential convergence, which concentrate onto the MMLE with quantitative nonasymptotic bounds on the estimation error ${\mathbb{E}}[\|\theta_{n}-\bar{\theta}_{\star}\|^{2}]^{1/2}$. In particular, we show that KIPLMC1 attains an accelerated error rate for error $\varepsilon$ after $\widetilde{\mathcal{O}}(\sqrt{d_{x}}\varepsilon^{-1})$ steps, compared to the IPLA and PGD which require $\widetilde{\mathcal{O}}(d_{x}\varepsilon^{-2})$ steps. The proposed algorithms both perform well empirically compared to the MPGD (with equal choice of step-sizes and friction coefficients) and other particle methods. Specifically, the induced noise should improve guarantees in the non-convex setting, as discussed in Akyildiz et al. [1], Raginsky et al. [44], Akyildiz and Sabanis [3].
The results presented here hold under assumptions of gradient-Lipschitzness and strong log-concavity, however, it is possible to extend this line of work to the non-convex, super-linear or even proximal cases (for examples of this see, Zhang et al. [53], Akyildiz and Sabanis [3], Johnston et al. [32], Cordero-Encinar et al. [17]). Another interesting direction is to consider the setting in Fan et al. [26], using an accelerated algorithm for high-dimensional linear models. Our methods can also be adapted to solving linear inverse problems as done by Akyildiz et al. [4], Glyn-Davies et al. [28]. Similar ideas can be considered within the setting of Stein variational gradient descent as done in Sharrock et al. [47]. Crucially, the procedure used to obtain the bounds for the algorithms presented here can easily be extended to other cases, creating a powerful theoretical framework for further underdamped Langevin sampling algorithms for the MMLE problem.
@@ Appendix A Preliminary results
In this section, we prove a key result that rewrites the KIPLD as a single underdamped Langevin diffusion evolving in ${\mathbb{R}}^{d_{\theta}+Nd_{x}}$ with a rescaled potential function and friction coefficient. This will be useful in the analysis of the KIPLD and its discretisation, as it allows us to leverage existing results on underdamped Langevin diffusions.
Proposition A.1.
Let
$(\bm{\theta}_{t},{\mathbf{X}}_{t}^{1},\ldots,{\mathbf{X}}_{t}^{N},{\mathbf{V}}_{t}^{\theta},{\mathbf{V}}_{t}^{x_{1}},\ldots,{\mathbf{V}}_{t}^{x_{N}})_{t\geq 0}$
solve (KIPLD). Define
$\widetilde{\bm{\theta}}_{t}$ $:=\bm{\theta}_{\sqrt{N}t},\qquad\widetilde{{\mathbf{X}}}_{t}^{i}:=N^{-1/2}{\mathbf{X}}^{i}_{\sqrt{N}t},$
$\widetilde{{\mathbf{V}}}_{t}^{\theta}$ $:=\sqrt{N}{\mathbf{V}}^{\theta}_{\sqrt{N}t},\qquad\widetilde{{\mathbf{V}}}_{t}^{x_{i}}:={\mathbf{V}}^{x_{i}}_{\sqrt{N}t},$
for $i\in[N]$, and set
$\widetilde{{\mathbf{Z}}}_{t}:=(\widetilde{\bm{\theta}}_{t},\widetilde{{\mathbf{X}}}_{t}^{1},\ldots,\widetilde{{\mathbf{X}}}_{t}^{N})^{\intercal}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}},\widetilde{{\mathbf{V}}}_{t}^{z}:=(\widetilde{{\mathbf{V}}}_{t}^{\theta},\widetilde{{\mathbf{V}}}_{t}^{x_{1}},\ldots,\widetilde{{\mathbf{V}}}_{t}^{x_{N}})^{\intercal}\in{\mathbb{R}}^{d_{\theta}+Nd_{x}},$
and
$\bar{U}_{N}(\theta,x_{1},\ldots,x_{N})$ $:=\frac{1}{N}\sum_{i=1}^{N}U(\theta,\sqrt{N}\,x_{i}),$ (11)
and $\widetilde{\gamma}:=\sqrt{N}\gamma$. Then there exists a ${\mathbb{R}}^{d_{\theta}+Nd_{x}}$-valued Brownian motion $(\widetilde{{\mathbf{B}}}_{t})_{t\geq 0}$ such that
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{z}$ $=-\widetilde{\gamma}\,\widetilde{{\mathbf{V}}}_{t}^{z}\,{\mathrm{d}}t-N\nabla_{z}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})\,{\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}.$
@@
By definition of the rescaled variables, we have
${\mathrm{d}}\widetilde{\bm{\theta}}_{t}$ $=\sqrt{N}\,{\mathrm{d}}\bm{\theta}_{\sqrt{N}t}=\sqrt{N}\,{\mathbf{V}}^{\theta}_{\sqrt{N}t}{\mathrm{d}}t=\widetilde{{\mathbf{V}}}_{t}^{\theta}{\mathrm{d}}t.$
Fix $i\in[N]$ and set $s(t):=\sqrt{N}t$. From (KIPLD), ${\mathrm{d}}{\mathbf{X}}_{t}^{i}={\mathbf{V}}_{t}^{x_{i}}{\mathrm{d}}t$, so
${\mathrm{d}}{\mathbf{X}}^{i}_{s(t)}={\mathbf{V}}^{x_{i}}_{s(t)}{\mathrm{d}}s(t)=\sqrt{N}\,{\mathbf{V}}^{x_{i}}_{\sqrt{N}t}{\mathrm{d}}t.$
Hence
${\mathrm{d}}\widetilde{{\mathbf{X}}}_{t}^{i}=N^{-1/2}\,{\mathrm{d}}{\mathbf{X}}^{i}_{\sqrt{N}t}=N^{-1/2}\,\sqrt{N}\,{\mathbf{V}}^{x_{i}}_{\sqrt{N}t}{\mathrm{d}}t=\widetilde{{\mathbf{V}}}_{t}^{x_{i}}{\mathrm{d}}t,$
and therefore ${\mathrm{d}}\widetilde{{\mathbf{Z}}}_{t}=\widetilde{{\mathbf{V}}}_{t}^{z}{\mathrm{d}}t$. Next, define $\widetilde{{\mathbf{B}}}_{t}^{j}:=N^{-1/4}{\mathbf{B}}_{\sqrt{N}t}^{j}$ for $j=0,\ldots,N$; then each $\widetilde{{\mathbf{B}}}_{t}^{j}$ is a Brownian motion and ${\mathrm{d}}{\mathbf{B}}_{\sqrt{N}t}^{j}=N^{1/4}{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}^{j}.$ Using (KIPLD), set $s(t):=\sqrt{N}t$. Then
${\mathrm{d}}\widetilde{{\mathbf{V}}}_{t}^{\theta}$ $=\sqrt{N}\,{\mathrm{d}}{\mathbf{V}}^{\theta}_{s(t)}$
@@
$=\sqrt{N}\left[\left(-\gamma{\mathbf{V}}^{\theta}_{s(t)}-\frac{1}{N}\!\sum_{i=1}^{N}\nabla_{\theta}U(\bm{\theta}_{s(t)},{\mathbf{X}}^{i}_{s(t)})\right){\mathrm{d}}s(t)+\sqrt{\frac{2\gamma}{N}}\,{\mathrm{d}}{\mathbf{B}}^{0}_{s(t)}\right]$
$=-N\gamma{\mathbf{V}}^{\theta}_{\sqrt{N}t}{\mathrm{d}}t-\sum_{i=1}^{N}\nabla_{\theta}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}^{0}$
$=-\widetilde{\gamma}\,\widetilde{{\mathbf{V}}}_{t}^{\theta}{\mathrm{d}}t-N\nabla_{\theta}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}^{0},$
where we used ${\mathrm{d}}s(t)=\sqrt{N}{\mathrm{d}}t$ and
$N\nabla_{\theta}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})=\sum_{i=1}^{N}\nabla_{\theta}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}).$
Fix $i\in[N]$. Using again $s(t)=\sqrt{N}t$ and (KIPLD),
${\mathrm{d}}\widetilde{V}_{t}^{x_{i}}$ $={\mathrm{d}}{\mathbf{V}}^{x_{i}}_{s(t)}$
@@
$=\left(-\gamma{\mathbf{V}}^{x_{i}}_{s(t)}-\nabla_{x}U(\bm{\theta}_{s(t)},{\mathbf{X}}^{i}_{s(t)})\right){\mathrm{d}}s(t)+\sqrt{2\gamma}\,{\mathrm{d}}{\mathbf{B}}^{i}_{s(t)}$
$=-\sqrt{N}\gamma{\mathbf{V}}^{x_{i}}_{\sqrt{N}t}{\mathrm{d}}t-\sqrt{N}\,\nabla_{x}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}){\mathrm{d}}t+\sqrt{2\gamma\sqrt{N}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}^{i}$
$=-\widetilde{\gamma}\,\widetilde{{\mathbf{V}}}_{t}^{x_{i}}{\mathrm{d}}t-N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t}){\mathrm{d}}t+\sqrt{2\widetilde{\gamma}}\,{\mathrm{d}}\widetilde{{\mathbf{B}}}_{t}^{i},$
where
$N\nabla_{z_{i}}\bar{U}_{N}(\widetilde{{\mathbf{Z}}}_{t})=\sqrt{N}\,\nabla_{x}U(\bm{\theta}_{\sqrt{N}t},{\mathbf{X}}^{i}_{\sqrt{N}t}).$
Stacking coordinates and defining $\widetilde{{\mathbf{B}}}_{t}=(\widetilde{{\mathbf{B}}}_{t}^{0},\widetilde{{\mathbf{B}}}_{t}^{1},\ldots,\widetilde{{\mathbf{B}}}_{t}^{N})^{\intercal}$ gives (12). $\Box$
Next we show that the stationary measure of the dynamics in (12) is given by a Gibbs measure with potential function $N\bar{U}_{N}$.
@@
The measure $\tilde{\pi}$ invariant for the dynamics in (12) is given as
@@
This result follows directly by observing the rescaling given in Proposition A.1 and the standard form of the stationary measure for underdamped Langevin diffusions. We now show that $\bar{U}_{N}$ is $\mu$-strongly convex under H1.